# Learning Unified User Quantized Tokenizers for User Representation

### v1

@@ Introduction
MSDP (Fu et al. 2023): Multi-scale Stochastic Distribution Prediction model for learning user behavioral sequence representation, which takes the prediction on user’s behaviors distribution over a period of time as the self-supervision signal.
One4all (Shin et al. 2021):One4all is a general-purpose representation learning through large-scale pre-training, while U-MLP is one of its extended user targeting model which adds MLP user decoder to generate targeted users.
FOUND (Dou et al. 2025):FOUND is a general user modeling method pretrained by self-supervised stage and user-text alignment stage, which enables zero-shot transfer and few-shot user targeting.
@@
MaskNet (Wang, She, and Zhang 2021): MaskNet introduces multiplicative operation into DNN ranking system by proposing instance-guided mask which performs element-wise product both on the feature embedding and feedforward layers.
Evaluation Metrics
For User Representation including future behavior and user profile prediction, linear probe representation analysis is conducted on all annotated datasets, which is evaluated by AUC (Area Under the ROC Curve (Bradley 1997)) and KS (Kolmogorov-Smirnov (Berger and Zhou 2014)) metrics. For Recommendation Task, we utilize AUC and HR (Hit Rate) to evaluate the performance of models.
| Feature | Backbone | AUC | HR@10 | HR@20 |
| User Profile | Masknet | 0.6476 | 0.2308 | 0.2968 |
| (+)Tokenizer | Masknet | 0.6527 | 0.2678 | 0.4913 |
Performance Analysis
The experimental results demonstrate that U2QT significantly outperforms baseline methods across downstream tasks in our designed benchmark. Our model achieves superior performance in AUC and KS in all domains and outperforms the second-best baseline 3% / 9% on average in AUC / KS. Moreover, we utilize the pretrained user tokenizer as supplementary feature in existing methods(e.g., One4all and MSDP) to verify the effectiveness. The result illustrates that these baselines gain huge improvement and outperform the original performance 1.1%/2.9% at AUC/KS for One4all and 0.7%/2.2% at AUC/KS for MSDP.
Ablation Study
In this section, we conduct an ablation study to evaluate the contribution of each heterogeneous data source. The experiment involves systematically removing one data domain at a time and then pre-training a corresponding tokenizer for evaluation on our benchmarks. The results, summarized in Table 3, lead to the following key conclusions:
| Model | Data Volume | Training Time | Storage |
| FOUND | 20,000,000 samples | about 7 hours | 55GB |
| U2QT | 20,000,000 samples | about 2 hours | 0.65GB |
Efficiency Analysis
Codebook Analysis
CodeBook Utilization
To delve into the inner workings of our proposed model, we analyze the utilization rate of the shared and domain-specific codebooks across the two quantization layers. The utilization rate is defined as the percentage of unique codes used during the training process, providing insights (van der Maaten and Hinton 2008) into how effectively the model leverages its representation capacity. The results are visualized in Figure 4.From the analysis, we draw several key observations:
i) Hierarchical Abstraction and Shared-Specific Disentanglement: The observed utilization patterns from our codebooks demonstrate that our hierarchical and modular design effectively organizes knowledge at different abstraction levels. The more compact usage of the shared codebook, particularly at the first layer, indicates that a small set of cross-domain patterns is captured and isolated from domain-specific ones. Meanwhile, the consistently higher utilization of the second-layer and domain-specific codebooks reflects the model’s ability to further refine and combine fundamental patterns into more diverse, fine-grained domain representations. This synergy between hierarchical abstraction and knowledge separation validates the design’s strength in representing both broad and nuanced information efficiently.
ii) The varying utilization rates across specific domains align with the inherent nature of the data. For instance, the ”Tabular” domain shows the lowest utilization. This is expected, as tabular data, consisting of numerical sequences, possesses less semantic richness and diversity compared to textual data from domains like ”App” or ”Bill”. Consequently, it can be effectively represented by a smaller, more specialized set of codes, confirming that our model’s utilization patterns realistically reflect the underlying data characteristics.
Multi-source Visualization
Conclusion
In this paper, we propose Unified User Tokenizer (U2QT), a novel framework for user representation. To unify the representation of heterogeneous user behaviors from various sources, we introduce Q-Former, which encodes multi-domain data into unified query vectors. For more efficient storage and computation, we design a multi-view RQ-VAE that compresses user representations into discrete tokens. We conduct comprehensive experiments across various downstream tasks to evaluate the effectiveness of our approach. The results consistently demonstrate that U2QT delivers strong representation capability, efficient storage and computation, and flexible cross-task generalization.

### v2

@@ Appendix A Baselines
MSDP (Fu et al., 2023): Multi-scale Stochastic Distribution Prediction model for learning user behavioral sequence representation, which takes the prediction on user’s behaviors distribution over a period of time as the self-supervision signal.
One4all (Shin et al., 2021):One4all is a general-purpose representation learning through large-scale pre-training, while U-MLP is one of its extended user targeting model which adds MLP user decoder to generate targeted users.
FOUND (Dou et al., 2025):FOUND is a general user modeling method pretrained by self-supervised stage and user-text alignment stage, which enables zero-shot transfer and few-shot user targeting.
@@
MaskNet (Wang et al., 2021): MaskNet introduces multiplicative operation into DNN ranking system by proposing instance-guided mask which performs element-wise product both on the feature embedding and feedforward layers.
Appendix B Evaluation Metrics
For User Representation including future behavior and user profile prediction, linear probe representation analysis is conducted on all annotated datasets, which is evaluated by AUC (Area Under the ROC Curve (Bradley, 1997)) and KS (Kolmogorov-Smirnov (Berger and Zhou, 2014)) metrics. For Recommendation Task, we utilize AUC and HR (Hit Rate) to evaluate the performance of models.
Appendix C Implementation Detail
In this section, we present the detailed implementation of our U2QT framework. We first describe the hyperparameters used in our model. First, we utilize the Qwen3-Embedding-0.6B222https://huggingface.co/Qwen/Qwen3-Embedding-0.6B to derive embeddings in 1024-dim for mutli-source data. The multi-view RQ-VAE module consists of a two-layer shared codebook with 256 entries of dimension 768, as well as six two-layer source-specific codebooks comprising 256 entries of dimension 128. The model is optimized using AdamW with a learning rate of 1e-3. All experiments are conducted on a cluster of sixteen PPU 810e processors, each with 100 GB of memory.